OpenStudy (thesammysaurus):

I need help with a couple of questions 1) Is the simplified form of 2^square root of 3 + 3^square root of 3 rational? Yes or No Question 2) Simplify Square root of 5 (10-4^square root of 2)

OpenStudy (anonymous):

1. yes

OpenStudy (anonymous):

does number 2 have any multiple choice question?

OpenStudy (thesammysaurus):

Yes they are: A) 14 B)\[15\sqrt{2}\] C)\[5\sqrt{2}-4\sqrt{10}\] D)None of the above @Sorcerer

OpenStudy (anonymous):

then D is ur answer

OpenStudy (thesammysaurus):

Thank you! But you actually got 1. wrong it was No. I just submitted it thats how I know. 2 was right, thank you for helping me @Sorcerer

OpenStudy (thesammysaurus):

Its fine, Im just going to redo it :) Ill give you a medal for helping anyways though @Sorcerer